# relsizeFont size

### v2

@@ 5.1 Supervised Learning
In supervised learning tasks, including image classification, text classification, object detection, and visual question answering, we aim to reduce computational costs by selecting a subset of informative and representative samples for training.
@@ 5.1.1 Experimental Setup
Baselines. We compare our method with the state-of-the-art sample selection methods, which are either applicable to all tasks or designed for a specific task. Baselines that are applicable to all tasks include: 1) Randomselection of samples, 2) Moderatecoreset (Xia et al., 2023), 3) CCS(Zheng et al., 2022), 4) $\mathbb{D}^{2}$runing(Maharana et al., 2024), and 5) GraphCut(Iyer et al., 2021). For image classification and text classification, the task-specific baselines include selecting the most difficult samples based on: 1) Entropy(Coleman et al., 2020), 2) Forgettingscore (Toneva et al., 2019), 3) EL2N(Paul et al., 2021), 4) AUM(Pleiss et al., 2020), and 5) Variance(Swayamdipta et al., 2020). For object detection, we include selection based on the AL-MDN uncertainty (Choi et al., 2021), which captures the detector’s overall uncertainty for an image. For visual question answering, we include selection based on the Instruction score (Wei et al., 2023), which evaluates an image-text pair based on image-text matching degree and text length.
@@
Implementation. For all tasks, we extract image embeddings using CLIP (Radford et al., 2021) and text embeddings using Sentence-BERT (Reimers & Gurevych, 2019) due to their demonstrated performance. For visual question answering, we concatenate the image and text embeddings for each sample. To measure training difficulty, we use AUM for image classification, Variance for text classification, AL-MDN uncertainty for object detection, and Instruction score for visual question answering. We ablate the different training difficulty metrics in Sec. LABEL:sub@sec:abl and observe that there is no significant performance difference among them. We also perform a grid search on the hyperparameters, such as $k$ in the $k$NN-graph construction (see Appendix LABEL:sub@appendix:hyper_setting for details).
5.1.2 Results
To cover a wide range of sampling rates, we select subsets that contain $1\%$, $2\%$, $5\%$, $10\%$, $20\%$, $50\%$, and $70\%$ of the entire training set. All the results are averaged over 5 random seeds. Table LABEL:sub@tab:dataset_sum_main shows the results on four datasets that cover all the four tasks. The full results are provided in Appendix LABEL:sub@appendix:eva_result.
Baselines that select the most difficult samples, such as AUM and Forgetting, perform well in high-rate settings. However, these methods fall behind in low-rate settings. This is due to their limited coverage of easy samples, which are crucial for model training when fewer samples are selected (Sorscher et al., 2022). Methods that prioritize sample coverage, such as CCS and $\mathbb{D}^{2}$ Pruning, address this issue and perform well in low-rate settings. However, in high-rate settings, they cannot accurately determine the most important samples that preserve global structure. This results in a lack of representativeness in the selected samples and suboptimal performance. In contrast, our method integrates both global and local metrics to better identify important samples and employs importance-biased blue noise sampling to ensure representativeness. Therefore, our method consistently performs better than baselines across all sampling rates and datasets.
In text classification and visual question answering, we observe that decreasing the sampling rate does not significantly affect performance. This is because we are fine-tuning pretrained models, which provide sufficient knowledge for these tasks. By providing high-quality samples, our method significantly accelerates fine-tuning with negligible performance loss. For example, as shown in Table LABEL:sub@tab:okvqa, we achieve a $10$x speedup by using only $10\%$ of the dataset to fine-tune MiniGPT-4, with only a $0.3\%$ drop in accuracy compared to using the entire dataset.
@@ 5.2 Active Learning
Active learning (Settles, 2009) aims to reduce annotation effort by selecting a set of informative and representative samples from an unlabeled pool. These samples are then labeled to train models.
@@ 5.2.1 Experimental Setup
Baselines. We include the baselines from Sec. LABEL:sub@subsec:supervised that are applicable to unlabeled datasets: 1) Random, 2) CCS, 3) $\mathbb{D}^{2}$runing, and 4) GraphCut. Additionally, we include Prototypicality (Sorscher et al., 2022) designed for unlabeled datasets. This method selects the most difficult samples based on the prototypicality score, which is defined as the distance between samples and their corresponding $k$-means cluster center. Difficult samples are those far from the center, as they tend to be more ambiguous than the samples closer to the center.
Implementation. In the active learning scenario, using a pretrained supervised model like CLIP for feature extraction is not suitable, because the domain of the unlabeled data may not be covered by its pretraining data. Therefore, we extract the image embeddings with a self-supervised model, SwAV (Caron et al., 2020). In the baselines and our method, the prototypicality score is utilized to measure training difficulty.
@@ 5.2.2 Results
We select unlabeled samples with rates of $1\%$, $2\%$, $5\%$, $10\%$, $20\%$, $50\%$, and $70\%$ and report the results averaged over 5 random seeds in Table LABEL:sub@tab:active. In low-rate settings, other baselines perform worse than random selection due to the absence of labels, indicating their reliance on labeled data to achieve optimal performance. In contrast, our method consistently performs better than random selection and other baseline methods. This is because structural entropy compensates for missing labels by capturing community structure in datasets.
@@ 5.3.1 Experimental Setup
Datasets and models. We use the datasets commonly used in continual learning, including Permuted MNIST, Split MNIST, Split CIFAR10, Split CIFAR100, and Split Tiny-ImageNet. Permuted MNIST splits MNIST (LeCun et al., 1998) into $10$ segments, where a fixed permutation of the pixel order is applied to all images in each segment to simulate different distributions. Thus, it contains $10$ classification tasks with samples from different distributions. The other four datasets split the image classes in MNIST, CIFAR10, CIFAR100, and Tiny-ImageNet (Le & Yang, 2015) into $5$, $5$, $20$, and $20$ segments, respectively, and each segment corresponds to a different classification task. In alignment with prior studies (Borsos et al., 2020; Hao et al., 2024), we use increasingly complex models as dataset complexity grows: a two-layer MLP for Permuted MNIST, a four-layer CNN for Split MNIST, ResNet-18 for Split CIFAR10, and ResNet-18 with multi-head output (Zenke et al., 2017) for Split CIFAR100 and Split Tiny-ImageNet. The models are trained sequentially on each task while maintaining a fixed-size replay memory that contains an equal number of samples for each previous task. During training, samples from both the replay memory and the current task are used, with the replay samples weighted by a hyperparameter that controls their influence on the current task. After completing each task, a subset of samples is selected from the current task to replace a portion of the replay memory. We report the average accuracy on all tasks.
Baselines. We include all baselines from Sec. LABEL:sub@subsec:supervised to select samples that update the replay memory. Additionally, we include three state-of-the-art selection methods designed for continual learning: 1) iCaRL(Rebuffi et al., 2017)that selects samples whose average embedding closely approximates the average embedding of all samples, 2) Greedy Coreset(Borsos et al., 2020)that formulates the selection as a bilevel optimization problem and greedily selects samples such that the model trained on them minimizes the loss across the entire dataset, and 3) BCSR(Hao et al., 2024)that formulates the selection as a bilevel optimization problem on the probability simplex over samples and introduces a regularizer to control the number of selected samples.
Implementation. Consistent with the implementation in Sec. LABEL:sub@subsec:supervised, we extract the image embeddings with CLIP and measure training difficulty with AUM. For all baselines and our method, we perform a grid search on the weight of replay samples during training to determine its optimal value (see Appendix LABEL:sub@appendix:exp_setting for details).
@@ 5.3.2 Results
Following Borsos et al. (2020), we test replay memory sizes of $100$ and $200$ and report the results averaged over 5 random seeds in Table LABEL:sub@tab:continual. Our method consistently achieves better performance than baselines across all memory sizes and datasets. This is because the combination of node-level structural entropy and training difficulty captures both the global structure of samples and the model training dynamics in retaining knowledge.
@@
Entropy 75.54 78.18 91.54 96.16 61.53 62.72 47.97 51.51 17.99 18.62
Forgetting 75.30 77.47 92.13 96.37 59.56 61.38 53.06 54.68 19.55 19.56
EL2N 75.57 77.48 88.54 94.87 57.79 58.34 46.47 48.11 17.95 17.98
AUM 75.54 77.47 90.38 95.79 58.32 58.06 46.14 47.20 18.24 17.89
Variance 75.67 77.42 91.90 94.96 58.69 57.77 52.83 54.00 19.32 19.50
@@ 5.4 Ablation Study
We conduct ablation studies on supervised learning using CIFAR10 to evaluate the impact and behavior of each module in our method.
Effect of modules. We ablate the two key modules in our method, node-level structural entropy (SE) and importance-biased blue noise sampling (BNS), by replacing them with alternatives. Without SE, we score samples based solely on training difficulty (TD). Without BNS, we either select samples with the highest scores (HS) or use the message passing (MP) method from $\mathbb{D}^{2}$ pruning. The hyperparameter controlling message weights in MP is searched as in the original paper (Maharana et al., 2024). Table LABEL:sub@tab:aba_module shows the results. Using TD and HS yields the lowest performance. Incorporating either SE, BNS, or MP largely improves performance, demonstrating their individual effectiveness. Combining SE with either BNS or MP further improves the performance, indicating that SE complements both methods in selecting informative and representative samples. Notably, BNS achieves comparable performance to MP without the need for tuning the additional hyperparameter, demonstrating its effectiveness.
Difficulty metric Sampling rate Avg.
70% 50% 20% 10% 5% 2% 1%
Identity 94.58 92.66 86.67 78.50 67.68 53.21 44.43 73.96
Forgetting 94.86 93.99 88.11 79.06 70.45 53.99 46.43 75.27
EL2N 94.61 93.74 88.45 79.93 69.52 54.01 45.15 75.06
AUM 95.01 94.50 88.31 80.24 69.82 54.78 45.25 75.42
Figure 3: Ablation of $k$ in the $k$NN-graph construction.
@@ 5.5 Qualitative Analysis
We visualize the selection results of different methods when selecting $2\%$ of the samples from CIFAR10 by projecting them onto a two-dimensional plane using t-SNE (van der Maaten & Hinton, 2008). Fig. LABEL:sub@fig:quality shows the results of AUM, $\mathbb{D}^{2}$ Pruning, and our method. The full results are provided in Appendix LABEL:sub@appendix:quali. Methods that select the most difficult samples, such as AUM, oversample near several class boundaries and undersample in several classes that are easier to classify (Fig. LABEL:sub@fig:quality(a)). Methods that prioritize sample coverage, such as $\mathbb{D}^{2}$ Pruning, achieve a better sample coverage but still miss critical samples near class boundaries (Fig. LABEL:sub@fig:quality(b)). This gap indicates that these methods may not effectively preserve the global structure of samples. Our method well covers the data distribution, providing a set of informative and representative samples for model training (Fig. LABEL:sub@fig:quality(c)).
@@ 6 Conclusion
In this paper, we present a structural-entropy-based sample selection method for efficient and effective learning. The key idea behind our method is to decompose graph-level structural entropy to a node-level global metric using the Shapley value. This global metric is combined with a local metric, training difficulty, for selecting informative and representative samples. The effectiveness of our method is validated by comprehensive experiments on three learning scenarios. Although our method has proven effective, future work on the following aspects is still promising. First, automating the hyperparameter selection based on data characteristics can reduce the computational costs of the grid search. Second, improving the support for multimodal data could strengthen its performance across a wider range of tasks, such as infographics VQA (Mathew et al., 2022) and building foundation models (Yang et al., 2024).
Appendix A Proof of Proposition LABEL:sub@prop:se
@@
Appendix B Dataset Statistics and Detailed Experimental Setting
B.1 Supervised Learning

### v3

@@ 5.1 Supervised Learning
In supervised learning tasks, we aim to reduce computational costs by selecting a subset of informative and representative samples for training.
@@ 5.1.1 Experimental Setup
Baselines. We compare our method with the state-of-the-art sample selection methods, which are either applicable to all tasks or designed for a specific task. Baselines that are applicable to all tasks include: 1) Randomselection of samples, 2) Moderatecoreset (Xia et al., 2023), 3) CCS(Zheng et al., 2022), 4) $\mathbb{D}^{2}$runing(Maharana et al., 2024), and 5) GraphCut(Iyer et al., 2021). For image classification and text classification, the task-specific baselines include selecting the most difficult samples based on: 1) Entropy(Coleman et al., 2020), 2) Forgettingscore (Toneva et al., 2019), 3) EL2N(Paul et al., 2021), 4) AUM(Pleiss et al., 2020), and 5) Variance(Swayamdipta et al., 2020). We also include two widely used baselines that prioritize diversity in sample selection, including $\bm{k}$-means (Xu et al., 2003), which selects the samples closest to $k$-means clustering centers, and $\bm{k}$-DPP (Kulesza & Taskar, 2011), which employs a determinantal point process to encourage diversity. For object detection, we include selection based on the AL-MDN uncertainty (Choi et al., 2021), which captures the detector’s overall uncertainty for an image. For visual question answering, we include selection based on the Instruction score (Wei et al., 2023), which evaluates an image-text pair based on image-text matching degree and text length.
@@
Implementation. For all tasks, we extract image embeddings using CLIP (Radford et al., 2021) and text embeddings using Sentence-BERT (Reimers & Gurevych, 2019) due to their demonstrated performance in capturing the semantic similarities. For visual question answering, we concatenate the image and text embeddings for each sample. To measure training difficulty, we use AUM for image classification, Variance for text classification, AL-MDN uncertainty for object detection, and Instruction score for visual question answering. We ablate the different training difficulty metrics in Appendix LABEL:sub@appendix:add_ablat and observe no significant performance difference among them. We also perform a grid search on the hyperparameters, such as $k$ in the $k$NN-graph construction (see Appendix LABEL:sub@appendix:hyper_setting for details).
@@
5.1.2 Results
To cover a wide range of sampling rates, we select subsets that contain $1\%$, $2\%$, $5\%$, $10\%$, $20\%$, $50\%$, and $70\%$ of the entire training set. All the results are averaged over 5 runs. Table LABEL:sub@tab:dataset_sum_main shows the results on four datasets that cover all the four tasks. The full results are provided in Appendix LABEL:sub@appendix:eva_result.
Baselines that select the most difficult samples, such as AUM and Forgetting, perform well in high-sampling-rate settings. However, these methods fall behind in low-sampling-rate settings. This is due to their limited coverage of easy samples, which are crucial for model training when fewer samples are selected (Sorscher et al., 2022). Methods that prioritize sample coverage, such as CCS and $\mathbb{D}^{2}$ Pruning, address this issue and perform well in low-sampling-rate settings. However, in high-sampling-rate settings, they cannot accurately determine the most important samples that preserve global structure. This results in a lack of representativeness in the selected samples and suboptimal performance. Methods that prioritize diversity are competitive in low-sampling-rate settings because they ensure the coverage of the dataset. However, in high-sampling-rate settings, they face challenges in balancing diversity with sample importance, leading to suboptimal performance. In contrast, our method integrates both global and local metrics to better identify important samples and employs importance-biased blue noise sampling to ensure representativeness. Therefore, our method consistently performs better than baselines across all sampling rates and datasets.
In text classification and visual question answering, we observe that decreasing the sampling rate does not significantly affect performance. This is because we are fine-tuning pretrained models, which provide sufficient knowledge for these tasks. By providing high-quality samples, our method significantly accelerates fine-tuning with negligible performance loss. For example, as shown in Table LABEL:sub@tab:okvqa, we achieve a $10$x speedup by using only $10\%$ of the dataset to fine-tune MiniGPT-4, with only a $0.3\%$ drop in accuracy compared to using the entire dataset. Specifically, our method reduces the fine-tuning time on a single Nvidia Tesla V100 GPU from approximately 30 minutes to 3 minutes, adding only a negligible selection overhead of 2 seconds.
@@ 5.2 Active Learning
Active learning (Settles, 2009) aims to reduce annotation effort by selecting a set of informative and representative samples from an unlabeled pool. These samples are then labeled to train models. The key difference between active learning and supervised learning lies in the absence of labels during sample selection, which makes the selection process more challenging.
@@ 5.2.1 Experimental Setup
Baselines. We include the baselines from Sec. LABEL:sub@subsec:supervised that are applicable to unlabeled datasets. Additionally, we include Prototypicality (Sorscher et al., 2022) designed for unlabeled datasets. This method selects the most difficult samples based on the prototypicality score, which is defined as the distance between samples and their corresponding $k$-means cluster center. Difficult samples are those far from the center, as they tend to be more ambiguous than the samples closer to the center.
Implementation. In the active learning scenario, using a pretrained supervised model like CLIP for feature extraction is not suitable, because the domain of the unlabeled data may not be covered by its pretraining data. Therefore, we extract the image embeddings with a self-supervised model, SwAV (Caron et al., 2020). The prototypicality score is utilized to measure training difficulty.
@@ 5.2.2 Results
We select unlabeled samples with rates of $1\%$, $2\%$, $5\%$, $10\%$, $20\%$, $50\%$, and $70\%$ and report the results averaged over 5 random seeds in Table LABEL:sub@tab:active. In low-sampling-rate settings, other baselines perform worse than random selection due to the absence of labels, indicating their reliance on labeled data to achieve optimal performance. In contrast, our method consistently performs better than random selection and other baseline methods. This is because structural entropy compensates for missing labels by capturing the community structure in the datasets.
@@ 5.3.1 Experimental Setup
Datasets and models. We use the datasets commonly used in continual learning, including Permuted MNIST, Split MNIST, Split CIFAR10, Split CIFAR100, and Split Tiny-ImageNet. Permuted MNIST splits MNIST (LeCun et al., 1998) into $10$ segments, where a fixed permutation of the pixel order is applied to all images in each segment to simulate different distributions. Thus, it contains $10$ classification tasks with samples from different distributions. The other four datasets split the image classes in MNIST, CIFAR10, CIFAR100, and Tiny-ImageNet (Le & Yang, 2015) into $5$, $5$, $20$, and $20$ segments, respectively, and each segment corresponds to a different classification task. In alignment with prior studies (Borsos et al., 2020; Hao et al., 2024), we use increasingly complex models as dataset complexity grows: a two-layer MLP for Permuted MNIST, a four-layer CNN for Split MNIST, ResNet-18 for Split CIFAR10, and ResNet-18 with multi-head output (Zenke et al., 2017) for Split CIFAR100 and Split Tiny-ImageNet. We conduct the experiments under three settings: class-incremental, task-incremental, and task-free (Wang et al., 2024). In the task-incremental setting, the models are trained sequentially for each task while maintaining a fixed-size replay memory that contains an equal number of samples for each previous task. During training, samples from both the replay memory and the current task are used, with the replay samples weighted by a hyperparameter that controls their influence. After each task is completed, samples are selected from the current task to replace a portion of the replay memory. We report the average accuracy on all tasks, with the model unaware of a sample’s task during testing. The other two settings are introduced in Appendix LABEL:sub@appendix:eva_result.
Baselines. We include baselines from Sec. LABEL:sub@subsec:supervised to select samples that update the replay memory. Additionally, we include nine widely used selection methods for continual learning: 1) $\bm{k}$-center (Sener & Savarese, 2018), 3) Gradient Matching (Campbell & Broderick, 2019), 5) FRCL (Titsias et al., 2020), 7) iCaRL(Rebuffi et al., 2017), 8) Greedy Coreset(Borsos et al., 2020), 9) BCSR(Hao et al., 2024), 10) DER++ (Buzzega et al., 2020), 11) GSS (Aljundi et al., 2019), and 12) GCR (Tiwari et al., 2022). Detailed introduction of these methods is provided in Appendix LABEL:sub@appendix:eva_result.
Implementation. We extract the image embeddings with CLIP and measure training difficulty with AUM. For all baselines and our method, we perform a grid search on the weight of replay samples to determine its optimal value (see Appendix LABEL:sub@appendix:exp_setting for details).
@@ 5.3.2 Results
$\bm{k}$-center 78.17 79.75 94.39 96.61 61.47 62.74 51.16 53.10 18.87 18.90
Gradient Matching 77.30 79.27 95.39 97.54 61.65 62.65 54.13 56.29 19.19 19.00
FRCL 77.33 79.21 94.48 97.10 61.67 62.93 51.40 54.28 18.86 19.01
@@
We test replay memory sizes of $50$, $100$, $200$, and $400$ and report the results averaged over 5 random seeds. Due to space limitations, Table LABEL:sub@tab:continual shows the results for replay memory sizes of $100$ and $200$ under the class-incremental setting, excluding baselines from Sec. LABEL:sub@subsec:supervised that are neither the best nor the runner-up in any dataset or memory size. The full results are provided in Appendix LABEL:sub@appendix:eva_result. Our method consistently achieves better performance than baselines across all memory sizes and datasets. This is because the combination of node-level structural entropy and training difficulty captures both the global structure of samples and the model training dynamics in retaining knowledge.
@@ 5.4 Ablation Study
We conduct ablation studies on supervised learning using CIFAR10 to evaluate the impact and behavior of each module in our method. This section presents results on the effect of modules. Additional ablation studies are provided in Appendix LABEL:sub@appendix:add_ablat.
Effect of modules. We ablate the two key modules in our method, node-level structural entropy (SE) and importance-biased blue noise sampling (BNS), by replacing them with alternatives. Without SE, we score samples based solely on training difficulty (TD). Without BNS, we either select samples with the highest scores (HS) or use the message passing (MP) method from $\mathbb{D}^{2}$ pruning. The hyperparameter controlling message weights in MP is determined through grid search as in the original paper (Maharana et al., 2024). Table LABEL:sub@tab:aba_module shows the results. Using TD and HS yields the lowest performance. Incorporating either SE, BNS, or MP largely improves performance, demonstrating their individual effectiveness. Combining SE with either BNS or MP further improves the performance, indicating that SE complements both methods in selecting informative and representative samples. Notably, BNS achieves comparable performance to MP without the need for the time-consuming grid search on the additional hyperparameter, demonstrating its effectiveness.
@@ 5.5 Qualitative Analysis
We visualize the selection results of different methods when selecting $2\%$ of the samples from CIFAR10 by projecting them onto a two-dimensional plane using t-SNE (van der Maaten & Hinton, 2008). Fig. LABEL:sub@fig:quality shows the results of AUM, $\mathbb{D}^{2}$ Pruning, and our method. Each point is a selected sample, and the colored contours indicate the high-density regions for each class. The full results are provided in Appendix LABEL:sub@appendix:quali. Methods that select the most difficult samples, such as AUM, oversample near several class boundaries and undersample in several classes that are easier to classify (Fig. LABEL:sub@fig:quality(a)). Methods that prioritize sample coverage, such as $\mathbb{D}^{2}$ Pruning, achieve a better sample coverage but still miss critical samples near class boundaries (Fig. LABEL:sub@fig:quality(b)). This gap indicates that these methods may not effectively preserve the global structure of the samples. Our method well covers the data distribution, providing a set of informative and representative samples for model training (Fig. LABEL:sub@fig:quality(c)).
@@ 6 Conclusion
In this paper, we present a structural-entropy-based sample selection method for efficient and effective learning. The key idea behind our method is to decompose graph-level structural entropy to a node-level global metric using the Shapley value. This global metric is combined with a local metric, training difficulty, for selecting informative and representative samples. The effectiveness of our method is validated by comprehensive experiments in three learning scenarios. Although our method has proven effective, future work on the following aspects is still promising. First, automating the hyperparameter selection based on data characteristics can reduce the computational costs of the grid search. Second, improving the support for multimodal data could strengthen its performance across a wider range of tasks, such as infographics VQA (Mathew et al., 2022) and building foundation models (Yang et al., 2024). Third, applying techniques such as hyper-class representation (Zhang et al., 2022) to enhance the image and text embeddings can improve the quality of selected samples.
7 Acknowledgements
Tianchi Xie, Jiangning Zhu, Minzhi Lin, and Shixia Liu are supported by the National Natural Science Foundation of China under grant U21A20469 and in part by the Tsinghua University-China Telecom Wanwei Joint Research Center. The authors would like to thank Duan Li, Yukai Guo, and Zhen Li for their valuable discussions and comments.
Appendix A Comparison of Encoding Tree Construction Methods
The construction of the encoding tree is an important step in our method. To the best of our knowledge, there are two methods for constructing the encoding tree: the one proposed by Li & Pan (2016) and the one proposed by Zhu et al. (2023). Both methods adopt the variations of the Huffman tree to construct the encoding tree. The key difference is that Zhu et al. (2023) further compresses the tree to a certain height to improve efficiency in subsequent processing. We assess the two methods by calculating the Pearson correlation of the resulting node-level structural entropy. As shown in Table LABEL:sub@tab:SE_constr, the correlations on CIFAR10, CIFAR100, and ImageNet-1K exceed $0.99$. Therefore, the choice of the construction method has a negligible effect on the sample selection results. In this paper, we employ the more recent method proposed by Zhu et al. (2023).
Table 5: Pearson correlation of the node-level structural entropy obtained from the two methods on CIFAR10, CIFAR100, and ImageNet-1K.
CIAFR10 CIFAR100 ImageNet-1K
Correlation 0.996 0.999 0.992
Appendix B Proof of Proposition LABEL:sub@prop:se
@@
Appendix C Theoretical Analysis of Node-Level Structural Entropy
C.1 Node-Level Structural Entropy and Sample Coverage
In this section, we prove that node-level structural entropy serves as a lower bound for the sample coverage. Following previous work (Zheng et al., 2022), we assume that the dataset $S=\{\mathbf{x}_{i},y_{i}\}$ contains $n$ i.i.d. samples drawn from an underlying distribution $P_{\mu}$ with probability measure $\mu$. We discuss a certain sample $u$’s ability to cover the entire distribution, i.e., how well the sample can cover the entire distribution. The sample coverage of a sample $\mathbf{u}$ can be quantified by the coverage of the entire distribution by a ball $B(\mathbf{u},r)$ centered at $\mathbf{u}$ with radius $r$ (Zheng et al., 2022):
$P(u,r)=\int_{B(u,r)}\text{d}\mu(x).$
When $r$ is fixed, the larger $P(u,r)$ is, the better sample covers the distribution. Directly calculating $P(u,r)$ is difficult because we do not know the underlying distribution $P_{\mu}$. We next prove that node-level structural entropy serves as a lower bound for the sample coverage.
We assume that the edge weights are in the range $[L,R]$, where $R>L>0$. Consider a node $u$ connected to $k$ other nodes $\{v_{1},v_{2},\dots,v_{k}\}$ with edge weights $\{w_{1},w_{2},\dots,w_{k}\}$. In the BFS tree rooted at $u$, let the total $vol$ of the subtrees of $v_{1},v_{2}\dots,v_{k}$ be $vol_{1},vol_{2},\dots,vol_{k}$. From the perspective of $u$, the calculation of $S_{e}(u)$ is as follows: $vol(u)$ is initially set to 0; For each node $v_{i}$ connected to $u$, add $vol_{i}$ of $v_{i}$ to $vol(u)$, and then add $w_{i}vol(u)$ to $S_{e}(u)$. If the nodes are connected in the oreder $\{\pi_{1},\pi_{2},\dots,\pi_{k}\}$, the results is:
$S_{e}(u)=\sum_{i=1}^{k}w_{\pi_{i}}\log(vol_{\pi_{1}}+vol_{\pi_{2}}+...vol_{\pi_{k}}).$
In practice, the nodes are connected in a specific order to minimize $S_{e}(u)$. Consequently, $S_{e}(u)$ is less than or equal to values obtained in any other order. Consider the order in which nodes are connected from the smallest $w_{i}$ to the largest $w_{i}$. We derive that:
$\begin{split}S_{e}(u)&\leq\sum_{i=1}^{k}w_{\pi_{i}}\log(vol_{\pi_{1}}+vol_{\pi_{2}}+...vol_{\pi_{k}})\\ &\leq\left(\sum_{i=1}^{k}w_{i}\right)\log\sum_{i=1}^{k}\left(\frac{w_{\pi_{k}}}{\sum_{i=1}^{k}w_{i}}(vol_{\pi_{1}}+...+vol_{\pi_{k}})\right)(\text{Jensen’s Inequality})\\ &=\left(\sum_{i=1}^{k}w_{i}\right)\log\sum_{i=1}^{k}(vol_{\pi_{k}}(w_{\pi_{1}}+w_{\pi_{2}}+...w_{\pi_{k}})/\sum_{i=1}^{k}w_{i})\\ &\leq kR\log\sum_{i=1}^{k}(vol_{\pi_{k}}(w_{\pi_{1}}+w_{\pi_{2}}+...w_{\pi_{k}})/\sum_{i=1}^{k}w_{i})\\ &\leq kR\log\sum_{i=1}^{k}(kvol_{\pi_{k}}w_{\pi_{k}}/\sum_{i=1}^{k}w_{i})\\ \end{split}$ (18)
Because $vol_{\pi_{k}}\in[kLn_{\pi_{k}},kRn_{\pi_{k}}]$, where $n_{\pi_{k}}$ is the number of nodes in the subtree of $v_{\pi_{k}}$. So we have:
$S_{e}(u)\leq kR\log(k^{2}R\sum_{i=1}^{k}n_{\pi_{i}}w_{\pi_{i}}/\sum_{i=1}^{k}w_{i})\$
Assume that all nodes in the subtree of $v_{i}$ are in $B(v_{i},r^{*})$. Then we have:
$\mathbb{E}[n_{i_{k}}]=n\int_{B(v_{i},r^{*})}\text{d}\mu=nP(v_{i},r^{*})$
which can then lead to:
$\begin{split}\mathbb{E}[S_{e}(u)]&\leq\mathbb{E}[kR\log\Big{(}k^{2}R(\sum_{i=1}^{k}n_{\pi_{i}}w_{\pi_{i}}/\sum_{i=1}^{k}w_{i})\Big{)}]\\ &\leq kR\log\mathbb{E}\Big{[}k^{2}R\left(\sum_{i=1}^{k}n_{\pi_{i}}w_{\pi_{i}}/\sum_{i=1}^{k}w_{i}\right)\Big{]}(\text{Jensen's Inequality})\\ &=kR\log\mathbb{E}\Big{[}nk^{2}R\left(\sum P(v_{\pi_{i}},r^{*})w_{\pi_{i}}/\sum_{i=1}^{k}w_{i}\right)\Big{]}\\ \end{split}$ (19)
Since
$\sum_{i=1}^{k}P(v_{\pi_{i}},r^{*})w_{\pi_{i}}/\sum_{i=1}^{k}w_{i}=\hat{P}(u,r)$
is actually the Nadaraya–Watson estimator of $P(u,r^{*})$, we have:
$\mathbb{E}[S_{e}(u)]\leq KR\log\mathbb{E}\Big{[}nK^{2}R\hat{P}(u,r)\Big{]}\$
Thus, the following theorem is obtained:
Theorem 1. For $r>r^{*}$, $S_{e}(u)$ ensures an lower-bound of the Nadaraya–Watson estimator $\hat{P}(u,r^{*})$ of $P(u,r^{*})$ with inequality:
$\frac{1}{nK^{2}R}\exp\left(\frac{1}{KR}\mathbb{E}[S_{e}(u)]\right)\leq\mathbb{E}[\hat{P}(u,r)]\approx\mathbb{E}[P(u,r)]\$
where $r^{*}$ is a constant related to the distribution $P_{\mu}$ and the location of $u$ in the metric space.
Theorem 1 shows the correlation between $S_{e}(u)$ and $P(u,r)$. For a certain sample $u$, a larger $S_{e}(u)$ guarantees a stronger sample coverage.
We further conduct an empirical validation to demonstrate that $\frac{1}{nk^{2}R}\exp(\frac{1}{kR}S_{e}(u))$ provides a relatively tight lower bound for $P(u,r)$. To precisely calculate $P(u,r)$, we simulate the data using a Gaussian Mixture Model, where the class centers $\mu_{1},\dots,\mu_{C}$ are drawn from a standard Gaussian distribution $\mathcal{N}(0,I)$. For each class $c$, the samples are drawn from $\mathcal{N}(\mu_{c},I)$. We align the synthetic data with real-world scales by generating 10 classes with 5,000 samples each, matching the statistics of CIFAR10. Our results show that: 1) For $90$% of the samples, $P(u,r)/\big{(}\frac{1}{nk^{2}R}\exp(\frac{1}{kR}S_{e}(u))\big{)}$ lies between $1.00$ and $1.45$; 2) For $99$% of the samples, $P(u,r)/\big{(}\frac{1}{nk^{2}R}\exp(\frac{1}{kR}S_{e}(u))\big{)}$ is less than $1.97$. These findings support the theorem that maximizing $S_{e}(u)$ during selection improves sample coverage.
C.2 Sample Coverage and Model Performance
Previous work (Zheng et al., 2022) has demonstrated that a better distribution coverage brings a lower empirical loss, as depicted by the following theorem:
Theorem 2. Zheng et al. (2022) Given $n$ i.i.d. samples drawn from $P_{\mu}$ as $S=\{\bm{x_{i}},y_{i}\}_{i\in[n]}$ where $\bm{x_{i}}\in X=\mathbb{R}^{d}$ is the feature representation of example $i$ and $y_{i}\in[C]$ is the class label for example $i$, a coreset $S^{\prime}$ which is a $p$-partial $r$-cover for $P_{\mu}$ on the feature space $X$, and an $\epsilon>1-p$, if the loss function $l(\cdot,y,w)$ is $\lambda_{l}$Lipschitz continuous for all $y$, $w$ and bounded by $L$, the class-specific regression function $\eta_{c}(x)=p(y=c|\bm{x})$ is $\lambda_{\eta}$-Lipschitz for all $c$, and $l(\bm{x},y;h_{S^{\prime}})=0$, $\forall(\bm{x},y)\in S^{\prime}$, then with probability at least $1-\epsilon$:
$\left|\frac{1}{n}\sum_{\bm{x},y\in S}l(x,y;h_{S^{\prime}})\right|\leq r(\lambda_{l}+\lambda_{\eta}LC)+L\sqrt{\frac{\log\frac{p}{p+\epsilon-1}}{2n}}$ (20)
Theorem 2 suggests that better sample coverage results in a tighter bound on the empirical loss for the entire set. Given the relationship between $S_{e}(u)$ and sample coverage, selecting samples with high node-level structural entropy effectively enhances model performance.
Appendix D Dataset Statistics and Detailed Experimental Setting
D.1 Supervised Learning